Tier1 Operations Report 2016-02-17

From GridPP Wiki
Jump to: navigation, search

RAL Tier1 Operations Report for 17th February 2016

Review of Issues during the week 10th to 17th February 2016.
  • The OPN link to CERN continues to be busy - although less so than during the previous fortnight.
Resolved Disk Server Issues
  • GDSS671 (AliceDisk - D1T0) failed on Saturday (13th Feb). Following tests and a disk being replaced it was returned to service on Monday (15th Feb).
Current operational status and issues
  • There is a problem seen by LHCb of a low but persistent rate of failure when copying the results of batch jobs to Castor. There is also a further problem that sometimes occurs when these (failed) writes are attempted to storage at other sites. A recent modification has improved, but not completed fixed this.
  • The intermittent, low-level, load-related packet loss seen over external connections is still being tracked. Likewise we have been working to understand some remaining low level of packet loss seen within a part of our Tier1 network.
Ongoing Disk Server Issues
  • None
Notable Changes made since the last meeting.
  • The Castor repack instance was updated from version to
  • Ongoing Condor updates (to version 8.4.4) on batches of worker nodes.
Declared in the GOC DB
Service Scheduled? Outage/At Risk Start End Duration Reason
Whole Site SCHEDULED WARNING 23/02/2016 09:00 23/02/2016 12:00 3 hours Warning on Site during an infrastucture upgrade.
Advanced warning for other interventions
The following items are being discussed and are still to be formally scheduled and announced.

Listing by category:

  • Databases:
    • Switch LFC/3D to new Database Infrastructure.
  • Castor:
    • Update SRMs to new version (includes updating to SL6).
    • Update to Castor version 2.1.15.
    • Migration of data from T10KC to T10KD tapes (affects Atlas & LHCb data).
  • Networking:
    • Replace the UKLight Router. Then upgrade the 'bypass' link to the RAL border routers to 2*10Gbit.
  • Fabric
    • Firmware updates on remaining EMC disk arrays (Castor, LFC)
  • Grid Services
    • A Load Balancer (HAProxy) will be used in front of the FTS service.
Entries in GOC DB starting since the last report.
  • None
Open GGUS Tickets (Snapshot during morning of meeting)
GGUS ID Level Urgency State Creation Last Update VO Subject
119389 Green Urgent Waiting Reply 2016-02-05 2016-02-15 LHCb Data transfer problem to RAL BUFFER
118809 Yellow Urgent On Hold 2016-01-05 2016-01-13 Towards a recommendation on how to configure memory limits for batch jobs
117683 Green Less Urgent In Progress 2015-11-18 2016-02-16 CASTOR at RAL not publishing GLUE 2
116864 Red Urgent In Progress 2015-10-12 2016-02-16 CMS T1_UK_RAL AAA opening and reading test failing again...
109358 Green Less Urgent Waiting Reply 2015-02-05 2016-02-09 SNO+ RAL WMS unavailable
Availability Report

Key: Atlas HC = Atlas HammerCloud (Queue ANALY_RAL_SL6, Template 508); CMS HC = CMS HammerCloud

Day OPS Alice Atlas CMS LHCb Atlas HC CMS HC Comment
10/02/16 100 100 100 100 100 97 100
11/02/16 100 100 100 100 100 100 N/A
12/02/16 100 100 100 100 100 94 100
13/02/16 100 100 100 100 100 100 N/A
14/02/16 100 100 100 100 100 100 100
15/02/16 100 100 97 100 100 93 100 Single SRM Test failure on Put (could not open connection to srm-atlas.gridpp.rl.ac.uk)
16/02/16 100 100 100 100 100 88 100